# Overcoming Reward Overoptimization via Adversarial Policy Optimization with Lightweight Uncertainty Estimation

### v1

@@ 3.2 Uncertainty via Last Layer Embeddings
Many methods for quantifying uncertainty in a neural network’s output, guided by the similarity derived from the model’s final layer embeddings, have demonstrated their effectiveness both in theory and in practice [44, 31]. In this work, we take a closer look at the following two methods due to their computational efficiency and theoretical soundness.
Uncertainty Quantification in Neural Bandits.
Previous work [44, 31] demonstrated that with probability $1-\delta$, the following inequality concerning the uncertainty, or the width of the confidence interval of the estimated reward, holds:
$|r^{*}(x,y)-r_{\varphi}(x,y)|\leq b\sqrt{e(x,y)^{\top}M_{D}^{-1}e(x,y)},$ (3)
where $b$ is a function of $\delta$ (typically the smaller $\delta$ is, the larger $b$ is). And $M_{D}$ summarizes all last layer embeddings observed in the training data for the reward model, i.e., $M_{D}=\sum_{i=1}^{N}\sum_{y\in\{y_{c}^{i},y_{r}^{i}\}}e(x_{i},y)e(x_{i},y)^{\top}$. For ease of illustration, we then denote $U_{x,y}^{CI}=\sqrt{e(x,y)^{\top}M_{D}^{-1}e_{(}x,y)}$. Intuitively, if the new prompt-response pair $(x,y)$ is similar to samples in the training data, applying the inverse of $M_{D}$ would result in a small uncertainty $U_{x,y}^{CI}$; otherwise, the uncertainty will be high.
Bayesian Uncertainty.
Bayesian uncertainty modelling, such as Bayesian Linear Regression (BLR) or Gaussian Processes (GP) [42], also offers an elegant method to quantify uncertainty in closed form.
For simplicity of illustration, we first consider BLR in this section and subsequently describe GPs [42]. In BLR, the projection weights $\phi$ are presumed to follow a Gaussian prior, expressed as $p(\phi)=\mathcal{N}(\phi|\mathbf{m}_{0},\mathbf{S}_{0})$, with $\mathbf{m}_{0}$ being the prior mean and $\mathbf{S}_{0}$ the prior covariance matrix.
Then the predictive posterior mean $\mu_{\text{post}}$ and variance $\sigma^{2}_{\text{post}}$ of the reward for a new datapoint $(x,y)$ in BLR can be expressed in closed form as follows:
$p(r|x,y,\mathcal{D})$ $=\mathcal{N}(r|\mu_{\text{post}},\sigma^{2}_{\text{post}})$ (4)
$\mu_{\text{post}}$ $=e(x,y)^{\top}\mathbf{m}_{D},$ (5)
$\sigma_{\text{post}}$ $=\sqrt{e(x,y)^{\top}\mathbf{S}_{D}^{-1}e(x,y)+\epsilon^{2}}.$ (6)
where $\epsilon$ is the aleatoric noise, $\mathbf{S}_{D}$ and $\mathbf{m}_{D}$ are the closed form posterior mean and variance of the projection weights $\phi$ calculated based on the reward training data $D$. Now that we have presented BLR, it should be noted for GPs, the mean and covariance can be derived analogously by replacing every $e(x,y)^{\top}e(x^{\prime},y^{\prime})$ computation with $k(e(x,y),e(x^{\prime},y^{\prime}))$, where $k$ is a kernel function. Due to space constraints, please refer to Appendix D for a comprehensive presentation of GPs and BLR.
Note that both aforementioned uncertainty estimation methods are Mahalanobis distances [23] in the embedding spaces, with different scaling covariances.
@@ 3.3 Empirical Effectiveness of Lightweight Uncertainty Estimation
We empirically investigate the effectiveness of the two lightweight uncertainty estimation methods introduced in Section 3.2. Since our objective is to harness uncertainty in reward modelling to mitigate the overoptimization issue in RLHF, we specifically examine whether the two approaches can identify discrepancies between estimated proxy rewards and gold rewards during the RL stage. We adopt a synthetic setup widely used in the literature [12, 9, 11], where we train a significantly larger “gold-standard” reward model that simulates human preferences and provides labels for training a proxy reward model.
We run a standard RLHF pipeline on the Anthropic HH [2] and TL;DR dataset[36], with reward and policy models established from LLama-7B [37]. The preference dataset is labelled by the gold reward model. In particular, for the Anthropic HH dataset, the gold reward model is trained from Vicuna-13B [27], and for the TL;DR dataset, the gold reward model is trained from LLama-13B [37]. Additional details, such as gold/proxy reward model training, PPO implementation, etc., can be found in Appendix A.1.
During the PPO training stage, we log the generated samples every 10 steps. Subsequently, we compute their gold reward, proxy reward, as well as reward uncertainties associated with the proxy reward. We compare the following uncertainty methods:
  • $\bullet$
    CI: The uncertainty quantified through Eq.(3);
  • $\bullet$
    GP: Gaussian Process uncertainty through Eq.(6);
  • $\bullet$
    ENS-7B: Ensemble of three LLama7B reward models;
  • $\bullet$
    ENS-3B: Ensemble of three 3B reward models based on OpenLLaMA3B_v2 [13]. Since there are no officially released 3B models for LLaMA, we adopt OpenLLaMA [13] which is an open-source smaller reproduction of Meta AI’s LLaMA, showcasing comparable performance.
The results are presented in Figure 1. In Figure 1b and 1d, we depict the correlation between derived uncertainty [^1] Uncertainties calculated by different methods have different scales, so we normalized them to (0,1) for ease of comparison. and the difference between gold reward and proxy reward (i.e., $|r^{*}(x,y)-r_{\varphi}(x,y)|$). Ideally, an effective uncertainty measure should increase as the difference between gold and proxy reward increases. In Figure 1a and Figure 1c, we illustrate the progression of proxy rewards (blue dashed line), gold rewards (blue solid line), and uncertainties calculated by two lightweight uncertainty estimation methods, along with ENS-3B, throughout the PPO optimization steps. The reward values are indexed on the left y-axis, while the uncertainty is indexed on the right y-axis. These figures provide insight into whether the calculated uncertainties can effectively recognize the over-optimization issue.
Our key insights are summarized as follows:
(a) Anthropic HH (Step)
(b) Anthropic HH (Reward Diff)
@@
Figure 1: Comparison among lightweight uncertainty estimations. In Figure 1a and 1c, the blue lines with shaded areas depict the reward dynamics concerning optimization steps in PPO, where the solid and dashed lines represent gold and proxy rewards, respectively. The lines with dots denote the results from different uncertainty estimation methods. The reward values are indexed on the left y-axis, while the uncertainty is indexed on the right y-axis. In Figure 1b and 1d, we plot the correlation between uncertainty and the difference between gold reward and proxy reward.
@@
$\bullet$ The lightweight uncertainty estimation methods are effective in signalling over-optimizaton. We observe from Figure 1b and 1d that, as the difference between gold reward and proxy reward increases, the uncertainty calculated by CI and GP also rises. This demonstrates CI and GP indeed carry information on when the proxy reward is drifting away from the ground-truth reward. Furthermore, in Figure 1a and 1c, it is evident that when there is a divergence between golden and proxy rewards, indicating over-optimization, the uncertainty calculated by CI and GP generally increases with the optimization steps. This suggests the potential to leverage them to address the over-optimization issue. CI appears to outperform GP since it consistently increases with the divergence of gold and proxy rewards, while GP tends to stabilize after roughly $500$ steps, as shown in Figure 1c.
@@
$\bullet$ The lightweight uncertainty estimation surpasses reward ensembles with comparable parameter sizes. Compared to the two lightweight uncertainty estimation methods, ENS-3B appears to be less correlated with the divergence between gold and proxy rewards, particularly on the TL;DR dataset [36]. Contrary to our lightweight uncertainty estimation methods, the uncertainty calculated by ENS-3B does not exhibit a monotonically increasing trend with the reward difference between gold and proxy rewards. This is likely due to the fact the smaller reward models, in this case, 3B models, are not able to capture the preference well, thus leading to worse predictions.
4 Utilizing Uncertainty to Mitigate Reward Over-Optimization
Given that the lightweight uncertainty estimation methods in Section 3 capture the reliability of estimated rewards and demonstrate potential in mitigating the overoptimization issue, this section devises an effective way to leverage them to improve policy optimization. For subsequent sections, we primarily adopt CI in Eq.(3) to quantify uncertainties in reward modelling, as it demonstrates strong effectiveness, as shown in Figure 1.
The reward uncertainties enable the construction of a confidence interval containing the golden reward with high probability, as defined in Eq.(3). This motivates us to formulate a distributionally robust optimization problem centred around the reward model’s prediction confidence interval for policy improvement. Moving from a naïve point estimate of reward to an interval estimate, this approach prevents the policy optimization from being misled by high rewards with high uncertainty, which is the main cause of overoptimization.
As the lightweight uncertainty estimation focuses solely on the last layer embeddings, the reward uncertainty implies uncertainty in the learnt projection weight in the reward model. Specifically, let $\phi^{*}\in\mathbb{R}^{d}$ and $\hat{\phi}\in\mathbb{R}^{d}$ denote the optimal and estimated projection weights for the last layer to predict rewards, respectively. We have that: $\exists B\in\mathbb{R}_{+},\|\phi^{*}-\hat{\phi}\|_{M_{D}}^{2}\leq B$ (Lemma C.4 [44]). Recall that $M_{D}$ summarizes all last layer embeddings observed during the training of the reward model and can thus be precomputed.
Within the confidence interval of learnt projection weight, AdvPO searches for reward predictions that are most pessimistic about the current policy and then improve the policy against such rewards. Since the gold reward is guaranteed to be covered by the confidence interval with high probability, using the most pessimistic reward (i.e., the lowest possible reward) naturally avoids over-optimization. Formally, this formulates a distributionally robust optimization problem,
$\max_{\pi_{\theta}}$ $\min_{\|\phi-\hat{\phi}\|^{2}_{M_{D}}\leq B}\mathbb{E}_{x,y\sim\pi_{\theta}(\cdot|x)}\left[r_{\phi}(x,y)\right]-\mathbb{E}_{x,y_{\rm ref}}\left[r_{\phi}(x,y_{\rm ref})\right]-\beta\mathbb{E}_{x,y\sim\pi_{\theta}(\cdot|x)}\left[\mathbb{D}_{\text{KL}}\left[\pi_{\theta}(y|x)\middle\|\pi_{\rm SFT}(y|x)\right]\right],$ (7)
Here, with a bit of abuse of notation, we use $r_{\phi}(\cdot)$ to denote the reward obtained when using the projection weight $\phi$, while keeping the representation encoder unchanged.
Furthermore, to better align our adversarial search with how the reward model is obtained in Eq. (1), we also incorporate a reference response through pairwise preference comparison in our objective function. The reference response can be any acceptable answer, such as an annotated good response from users or a response generated by the SFT model. Intuitively, the inclusion of a reference also prevents AdvPO from being overly conservative, as it ensures that the model is optimized along the direction that can at least yield a policy better than where the reference is collected.
Theorem C.1 in Appendix C demonstrates the inner minimization of Eq.(7) has a closed-form solution, and thus the optimization problem in Eq.(7) has an equivalent but easier to operate form:
$\max_{\pi_{\theta}}\quad$ $\mathbb{E}_{x,y\sim\pi_{\theta}(\cdot|x)}[r_{\hat{\phi}}(x,y)-\frac{1}{\lambda^{*}}e(x,y)^{\top}M_{D}^{-1}g]-\mathbb{E}_{x,y_{\rm ref}}[r_{\hat{\phi}}(x,y_{\rm ref})-\frac{1}{\lambda^{*}}e(x,y_{\rm ref})^{\top}M_{D}^{-1}g],$
where $g=\mathbb{E}_{x,y\sim\pi_{\theta}(\cdot|x)}\left[e(x,y)\right]-\mathbb{E}_{x,y_{\rm ref}}\left[e(x,y_{\rm ref})\right]$ and $\lambda^{*}=\sqrt{\frac{g^{\top}M^{-1}g}{B}}$.
Comparison to previous approaches in utilizing uncertainty against overoptimization. Previous work [9, 11] utilizes reward uncertainty on a per-sample basis, i.e., penalizing each sample’s reward based on its individual uncertainty, as illustrated in Eq. (2). While both per-sample uncertainty penalization and AdvPO adopt a conservative approach to leveraging reward uncertainty, the degree of conservatism is crucial [16, 30]. Excessive conservatism, i.e., penalizing rewards too heavily based on uncertainties, is known to impede the discovery of the correct direction for optimization, thus failing to find a good policy.
We theoretically prove in Lemma C.2 in Appendix C that, compared to [9, 11], with same confidence interval on reward, AdvPO leverages uncertainty less conservatively, making it more likely to contribute to a better policy while mitigating overoptimization.
5 Experiments
In this section, we empirically assess AdvPO’s effectiveness. Our experiments focus on evaluating : (1) whether AdvPO can mitigate the overoptimization issue; (2) whether AdvPO results in an improved policy in practice.
Datasets. We utilized the following two widely adopted datasets for RLHF to carry out our empirical investigation.
  • $\bullet$
  • $\bullet$
5.1 AdvPO mitigates the over-optimization
To assess whether AdvPO can mitigate reward over-optimization, we perform experiments in the synthetic setup outlined in Section 3.3, where a significantly larger ”gold-standard” reward model is employed to simulate human preferences. This setup is consistent with the evaluation approach used in previous studies [9, 11]. Additional implementation details regarding gold reward training, proxy reward training dataset generation, proxy rewards model training, and RL optimization can be found in Appendix A.1.

### v2

@@ 3.2 Uncertainty via Last Layer Embeddings
Many methods, derived from a neural network model’s final layer embeddings, have demonstrated their effectiveness for quantifying uncertainty in the network predictions, both theoretically and in practice [43, 31]. In this work, we specifically follow the line of uncertainty quantification methods in neural bandits [43], due to its computational efficiency and theoretical soundness.
We first present the following theorem on reward uncertainties when learning rewards through the Bradley-Terry preference model, assuming the reward model architecture is infinitely wide.
Theorem 3.1.
Assuming the network architecture is infinitely wide and its neural tangent kernel matrix is positive definite, learning rewards through the Bradley-Terry preference model yields the following inequality for the width of the confidence interval of the estimated reward $r_{\hat{\varphi}}(x,y)$. Specifically, with probability $1-\delta$:
$|r^{*}(x,y)-r_{\hat{\varphi}}(x,y)|\leq b\sqrt{e(x,y)^{\top}M_{D}^{-1}e(x,y)}+\text{const},$ (4)
where $b$ is a term related to the preference dataset $D$ and $\delta$ (typically the smaller $\delta$ is, the larger $b$ is), $r^{*}$ and $r_{\hat{\varphi}}$ denote the unknown ground-truth reward and estimated reward model parameterized by $\hat{\varphi}$ respectively, and $M_{D}$ summarizes all last layer embeddings observed in the preference dataset $D$ for the reward model, i.e., $M_{D}=\lambda I+\sum_{i=1}^{N}\sum_{y\in\{y_{c}^{i},y_{r}^{i}\}}e(x_{i},y)e(x_{i},y)^{\top}$. Here $\lambda$ is a regulariser for the inversion.
Intuitively, Theorem 3.1 bounds the absolute difference between the predicted reward $r_{\hat{\varphi}}$ and the ground-truth reward $r^{*}$. Consequently, it is natural to define the uncertainty around the predicted reward $r_{\hat{\varphi}}(x,y)$ as
$U_{x,y}^{CI}=b\sqrt{e(x,y)^{\top}M_{D}^{-1}e(x,y)},$
since a larger difference between $r_{\hat{\varphi}}$ and $r^{*}$ implies greater uncertainty. This becomes even clearer when taking a closer look at $U_{x,y}^{CI}$. When a new prompt-response pair $(x,y)$ is similar to samples in the training data, applying the inverse of $M_{D}$, which is constructed using the training data, results in a smaller uncertainty $U_{x,y}^{CI}$. Conversely, if the pair diverges significantly from the training samples, the uncertainty $U_{x,y}^{CI}$ will be high. Note that the second term in Eq.(4) is constant and can thus be omitted when comparing reward uncertainties across prompt-response pairs. Finally, it is worth pointing out that computing $U_{x,y}^{CI}$ is computationally cheap at the policy training stage (i.e., $\mathcal{O}(d^{2})$, where $d$ is the dimension of the embeddings) as $M_{D}^{-1}$ can be pre-computed.
Remark on Assumptions in Theorem 3.1. The derivation of Eq. (4) relies on certain assumptions regarding network architectures, specifically that the network width is infinitely wide and neural tangent kernel matrix is positive definite. Recent work [24] that studied the Neural Tangent Kernel (NTK) in language models has also adopted similar assumptions, and its effectiveness suggests the reasonableness of these assumptions in LLMs.
Empirical verification. In Section 5.1, we empirically examine the effectiveness of the proposed lightweight uncertainty estimation using a synthetic setup with known ground-truth rewards. Our findings indicate that $U_{x,y}^{CI}$ accurately captures the divergence between the ground-truth and estimated proxy rewards, effectively signalling overoptimization.
Having detailed how to obtain uncertainty regions around the predicted reward, we will now illustrate in the next section how these uncertainty estimates can be used in policy optimization.
4 Utilizing Uncertainty to Mitigate Reward Overoptimization
This section introduces our framework, AdvPO, to address the issue of overoptimization during policy optimization by leveraging the aforementioned lightweight uncertainty estimation.
Instead of optimizing towards a potentially incorrect point estimate $r_{\hat{\varphi}}$, which causes overoptimization, AdvPO aim to optimize the following MaxMin objective which takes into account the confidence region around the imperfect reward model $r_{\hat{\varphi}}$ that contains the ground-truth reward $r^{*}$:
$\max_{\pi_{\theta}}$ $\min_{{\varphi\in C_{\delta}^{r}(\hat{\varphi})}}\mathbb{E}_{x,y\sim\pi_{\theta}(\cdot|x)}\left[r_{\varphi}(x,y)\right]-\beta\mathbb{E}_{x,y\sim\pi_{\theta}(\cdot|x)}\left[\mathbb{D}_{\text{KL}}\left[\pi_{\theta}(y|x)\middle\|\pi_{\rm SFT}(y|x)\right]\right],$
Here, $C_{\delta}^{r}(\hat{\varphi})$ contains all possible $\varphi$ values centered around the current estimate $\hat{\varphi}$, but also includes the optimal $\varphi^{*}$ that yields the ground truth reward, with a probability of $1-\delta$. Intuitively, AdvPO aims to maximize the same objective as in standard PPO (Eq. 2), while also adversarially searching for the pessimistic reward function within the predicted reward $r_{\hat{\varphi}}$’s confidence ball containing the ground-truth reward $r^{*}$. However, this MaxMin objective poses some practical issues, given the inner minimization over $C_{\delta}^{r}(\hat{\varphi})$ is intractable. Hence AdvPO makes the following observation.
Connection between Rewards and Projection Weights: An important corollary to Theorem 3.1, crucial to AdvPO, is that $U_{x,y}^{CI}$, the first term of the upper bound of the reward difference $|r^{*}(x,y)-r_{\varphi}(x,y)|$ in Eq.(4), actually represents the uncertainty stemming from the inaccuracy in the estimated projection weight $\hat{\phi}$.
Recall that $e(x,y)$ denotes the last layer embedding of the prompt and response pair $(x,y)$. Let $\phi^{*}$ and $\hat{\phi}$ be the optimal and estimated projection weights for reward prediction respectively. Under the assumption mentioned in Section 3.2, the ground-truth reward can be approximated by a linear function of optimal projection weight $\phi^{*}$ and $e(x,y)$, plus a term that can be bounded, i.e., $r^{*}(x,y)=e(x,y)^{\top}\phi^{*}+\text{bounded term}$. Considering the linearity of rewards with respect to the last layer embeddings $r_{\hat{\varphi}}(x,y)=\hat{\phi}^{\top}e(x,y)$, and denoting the established confidence region of the projection weight as $\|\hat{\phi}-\phi^{*}\|_{M_{D}}\leq b$, we show that:
$|r^{*}(x,y)-r_{\hat{\varphi}}(x,y)|\leq\underbrace{|e(x,y)^{\top}\phi^{*}-e(x,y)^{\top}\hat{\phi}|+\text{const}}_{A_{1}}\leq\underbrace{\|\phi^{*}-\hat{\phi}\|_{M_{D}}\sqrt{e(x,y)^{\top}M_{D}^{-1}e(x,y)}}_{\leq U_{x,y}^{CI}}+\text{const}$
Therefore, the objective of the inner optimization problem in AdvPO can be relaxed to optimize an upper bound, i.e., $A_{1}$ in Eq. (4), where the minimization is now taken over the projection weights instead of the reward functions.
$\max_{\pi_{\theta}}$ $\min_{{\|\phi-\hat{\phi}\|_{M_{D}}\leq b}}\mathbb{E}_{x,y\sim\pi_{\theta}(\cdot|x)}\left[r_{\phi}(x,y)\right]-\beta\mathbb{E}_{x,y\sim\pi_{\theta}(\cdot|x)}\left[\mathbb{D}_{\text{KL}}\left[\pi_{\theta}(y|x)\middle\|\pi_{\rm SFT}(y|x)\right]\right],$ (5)
Here, with a bit of abuse of notations, we use $r_{\phi}(\cdot)$ to denote the reward obtained when using the projection weight $\phi$, while keeping the representation encoder unchanged.
It is important to note that this approach diverges significantly from conventional reward penalization methods [9, 11, 45]. Rather than focusing on the worst-case scenario for each sample, our objective function adopts a more holistic perspective by minimizing across the reward functions themselves. Further details on the differences will be elaborated later in this section (Lemma 4.2).
Incorporating Reference Responses. To prevent AdvPO from becoming overly pessimistic, we introduce reference responses $\{y_{\rm ref}\}$ into the objective, resulting in the final objective of AdvPO:
[…]
$\quad-\beta\mathbb{E}_{x,y\sim\pi_{\theta}(\cdot|x)}\left[\mathbb{D}_{\text{KL}}\left[\pi_{\theta}(y|x)\middle\|\pi_{\rm SFT}(y|x)\right]\right],$ (6)
As illustrated in Lemma C.1, incorporating reference responses enforces policy optimization towards the direction of the reference responses, i.e., $\mathbb{E}_{x,y_{\rm ref}}[e(x,y_{\rm ref})]$, while optimizing against pessimistic rewards. Thus as long as the set of reference responses is reasonably good and achieves a positive ground-truth reward on average, i.e, $(\mathbb{E}_{x,y_{\rm ref}}[e(x,y_{\rm ref})])^{\top}\phi^{*}>0$, the policy is guided to outperform the reference, preventing AdvPO from being overly pessimistic. In practice, the reference responses can be any acceptable answers, such as annotated good responses from users or responses generated by the SFT policy.
Next, we show in Theorem 4.1 that the inner minimization of Eq.(6) has a closed-form solution and hence Eq.(6) reduces to an objective function that can be optimized using standard gradient ascent.
Theorem 4.1.
The optimization problem in Eq.(6) is equivalent to the following objective:
$\max_{\pi_{\theta}}\quad$ $\mathbb{E}_{x,y\sim\pi_{\theta}(\cdot|x)}[r_{\hat{\phi}}(x,y)-\frac{1}{\lambda^{*}}e(x,y)^{\top}M_{D}^{-1}g]-\mathbb{E}_{x,y_{\rm ref}}[r_{\hat{\phi}}(x,y_{\rm ref})-\frac{1}{\lambda^{*}}e(x,y_{\rm ref})^{\top}M_{D}^{-1}g]$
where $e(x,y)$ denotes the last layer embedding of the prompt-response pair $(x,y)$, and $g=\mathbb{E}_{x,y\sim\pi_{\theta}(\cdot|x)}\left[e(x,y)\right]-\mathbb{E}_{x,y_{\rm ref}}\left[e(x,y_{\rm ref})\right]$ and $\lambda^{*}=\sqrt{\frac{g^{\top}M^{-1}g}{b^{2}}}$.
Theorem 4.1 shows that the initial MaxMin objective can be loosened and rewritten into a standard Max objective for which we can use standard gradient ascent.
Comparison to previous approaches in utilizing uncertainty against overoptimization. As mentioned above, recent works [9, 11, 45] utilize reward uncertainty on a per-sample basis, i.e., penalizing each sample’s reward based on its individual uncertainty, as illustrated in Eq.(3). While both per-sample uncertainty penalization and AdvPO adopt a pessimistic approach to leverage reward uncertainty, the degree of pessimism is crucial [16, 30, 45]. Excessive pessimism, i.e., penalizing rewards too heavily based on uncertainties, is known to impede the discovery of the correct direction for optimization, thus failing to find a good policy. To demonstrate this, we prove the following:
Lemma 4.2.
This demonstrates that AdvPO is more effective in enhancing policy performance while reducing over-optimization, which we will back up with extensive large-scale experiments in the next section.
5 Experiments
In this section, we present our empirical results. In Section 5.1, we evaluate the effectiveness of the proposed lightweight uncertainty estimation. The effectiveness of AdvPO is demonstrated through (1) assessing whether ADVPO can mitigate the over-optimization issue in Section 5.1, and (2) examining whether AdvPO leads to an improved policy in practice in Section 5.3.
Datasets. We used two widely adopted datasets, Anthropic HH [2] and TL;DR [35], for empirical investigation. Additional dataset descriptions can be found in Appendix A.1.
@@ 5.1 Empirical effectiveness of lightweight uncertainty estimation
While our goal is to signal potential overoptimization during the RL stage, we specifically examine whether the quantified uncertainties $U_{x,y}^{CI}$ in Section 3.2 can identify discrepancies between estimated proxy rewards and ground-truth rewards during the RL stage. We adopt a synthetic setup widely used in the literature [12, 9, 11], where we train a significantly larger “gold-standard” reward model that simulates human preferences and provides labels for training a proxy reward model.
For both datasets, we trained a gold reward model using the LLama-13B model and established the reward and policy model in RLHF from the LLama-7B [36]. More details, such as gold/proxy reward model training, PPO implementation, etc., can be found in Appendix A.3.
We log the generated samples every 10 steps during the PPO training stage. Subsequently, we compute their gold reward, proxy reward, as well as reward uncertainties associated with the proxy reward. In addition to the lightweight uncertainty estimation methods (denoted as CI), we also investigate two ensemble-based uncertainty quantification methods: (1) ENS-7B: Ensemble of three LLama7B reward models; (2) ENS-3B: Ensemble of three 3B reward models based on OpenLLaMA3B_v2 [13], aiming to match the ensemble model size roughly comparable to CI, which quantifies uncertainties based on LLama7B. We adopt OpenLLaMA [13] as there are no official 3B LLama models. OpenLLaMA is an open-source smaller reproduction of Meta AI’s LLaMA, demonstrating comparable performance.
Note that ENS-7B has only been added for completeness. ENS-7B requires significantly more training, memory and inference compute compared to our proposed CI. Nevertheless, we believe that this side-by-side comparison illustrates the effectiveness of our lightweight uncertainty estimation.
Results. The results are presented in Figure 1. We have the following two key observations:
(a) Ant.HH (Step)
(b) Ant.HH (Reward Diff)
@@
Figure 1: Comparison among lightweight uncertainty estimations. In Figure 1a and 1c, the blue lines with shaded areas depict the reward dynamics concerning optimization steps in PPO, where the solid and dashed lines represent gold and proxy rewards, respectively. The lines with dots denote the results from different uncertainty estimation methods. The reward values are indexed on the left y-axis, while the uncertainty is indexed on the right y-axis. In Figure 1b and 1d, we plot the correlation between uncertainty and the difference between gold and proxy rewards.
@@
$\bullet$ The lightweight uncertainty effectively captures the discrepancy between proxy and gold rewards, signalling over-optimization. First, we observe from Figure 1b and 1d that, as the difference between gold and proxy rewards increases, the uncertainty calculated by our CI also rises. This demonstrates that our proposed CI indeed captures information about when the proxy reward is drifting away from the ground-truth reward. Furthermore, in Figure 1a and 1c, it is evident that when there is a divergence between proxy rewards (blue dashed line) and gold rewards (blue solid line), indicating overoptimization, the uncertainty calculated by CI (red line) generally increases with the optimization steps. This suggests the potential to leverage them to address the overoptimization issue.
@@
$\bullet$ The lightweight uncertainty estimation surpasses reward ensembles with comparable parameter sizes. Compared to CI, ENS-3B appears to be less correlated with the divergence between gold and proxy rewards, particularly on the TL;DR dataset [35]. As shown in Figure 1c and 1d, unlike our method CI (red line), the uncertainty calculated by ENS-3B (grey line) does not exhibit a monotonically increasing trend with the reward difference between gold and proxy rewards. This is likely due to the fact the smaller reward models, in this case, 3B models, are not able to capture the preference well, thus leading to worse predictions.
As an ablation, we tested out a configuration using five 3B reward models in Appendix LABEL:appendix-subsec-ens-3b-5. However, we observe that increasing the number of models in the ensemble does not mitigate the failures observed with ENS-3B on TL;DR datasets. Moreover, for completeness, we also investigated the effectiveness of other uncertainty quantification methods such as using Bayesian uncertainty on last-layer embeddings. For more details and experimental results, we refer to Appendix D.
5.2 AdvPO mitigates reward overoptimization
Next, we transition to evaluating the effectiveness of AdvPO. We begin by assessing whether AdvPO can mitigate reward over-optimization under the same synthetic setup described in Section 5.1, where a significantly larger “gold-standard” reward model is used to simulate human preferences.